# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ Appendix E Reinforcement Learning Generated Subgraph Details
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0220\scriptstyle\pm 0.0249 0.0$$367\scriptstyle\pm 0.0404 0.0$$275\scriptstyle\pm 0.0308 0.0$$141\scriptstyle\pm 0.0159 0.0$$\scriptstyle\pm 0.0000 0.0000±0.00$$0\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $0375\scriptstyle\pm 0.0385 0.0$$067\scriptstyle\pm 0.0058 0.0$$111\scriptstyle\pm 0.0097 0.0$$056\scriptstyle\pm 0.0049 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(QA) $0.3985\scriptstyle\pm 0.0669$ $0.3900\scriptstyle\pm 0.0721$ $0.3942\scriptstyle\pm 0.0695$ $0.2470\scriptstyle\pm 0.0530$ $0.9333\scriptstyle\pm 0.1155$ $\mathbf{0.8333}\scriptstyle\pm\mathbf{0.0577}$
L3-FT(QA) + KG $0.4200\scriptstyle\pm 0.0230$ $0.4100\scriptstyle\pm 0.0173$ $0.4149\scriptstyle\pm 0.0200$ $0.2619\scriptstyle\pm 0.0158$ $0.8667\scriptstyle\pm 0.1155$ $0.7000\scriptstyle\pm 0.1000$
G-Retriever $0.3196\scriptstyle\pm 0.0307$ $0.2767\scriptstyle\pm 0.0058$ $0.2963\scriptstyle\pm 0.0168$ $0.1740\scriptstyle\pm 0.0115$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8000\scriptstyle\pm 0.0000$
GALAX $\mathbf{0.4400}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.4400}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.4400}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.2821}\scriptstyle\pm\mathbf{0.0000}$ $0.8667\scriptstyle\pm 0.1155$ $0.6333\scriptstyle\pm 0.2309$
Table 30: Model performances on $M$
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0275\scriptstyle\pm 0.0296 0.0$$167\scriptstyle\pm 0.0208 0.0$$156\scriptstyle\pm 0.0149 0.0$$079\scriptstyle\pm 0.0075 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(Med) + KG $1091\scriptstyle\pm 0.0630 0.0$$200\scriptstyle\pm 0.0000 0.0$$329\scriptstyle\pm 0.0027 0.0$$0167\scriptstyle\pm 0.0014$ $0.1333\scriptstyle\pm 0.1155$ $0.1333\scriptstyle\pm 0.1155$
L3-FT(QA) $0.4001\scriptstyle\pm 0.0022$ $0.3800\scriptstyle\pm 0.0265$ $0.3894\scriptstyle\pm 0.0134$ $0.2418\scriptstyle\pm 0.0103$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.0000$
L3-FT(QA) + KG $0.4753\scriptstyle\pm 0.0613$ $0.4367\scriptstyle\pm 0.0306$ $0.4549\scriptstyle\pm 0.0447$ $0.2952\scriptstyle\pm 0.0377$ $0.8000\scriptstyle\pm 0.0000$ $0.7333\scriptstyle\pm 0.0577$
G-Retriever $0.4514\scriptstyle\pm 0.0890$ $0.3667\scriptstyle\pm 0.0808$ $0.4046\scriptstyle\pm 0.0850$ $0.2561\scriptstyle\pm 0.0690$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.0000$
@@
Table 31: Model performances on KIRC
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5458\scriptstyle\pm 0.0450$ $0.5333\scriptstyle\pm 0.0513$ $0.5394\scriptstyle\pm 0.0478$ $0.3703\scriptstyle\pm 0.0452$ $0.8667\scriptstyle\pm 0.2309$ $0.8667\scriptstyle\pm 0.1155$
L3-FT(QA) + KG $0.4897\scriptstyle\pm 0.0182$ $0.4800\scriptstyle\pm 0.0265$ $0.4847\scriptstyle\pm 0.0219$ $0.3201\scriptstyle\pm 0.0192$ $0.8667\scriptstyle\pm 0.1155$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.0000}$
G-Retriever $0.4160\scriptstyle\pm 0.0315$ $0.3733\scriptstyle\pm 0.0058$ $0.3933\scriptstyle\pm 0.0171$ $0.2449\scriptstyle\pm 0.0133$ $0.6000\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$
GALAX $\mathbf{0.5797}\scriptstyle\pm\mathbf{0.0251}$ $\mathbf{0.5700}\scriptstyle\pm\mathbf{0.0265}$ $\mathbf{0.5748}\scriptstyle\pm\mathbf{0.0254}$ $\mathbf{0.4036}\scriptstyle\pm\mathbf{0.0247}$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.1155}$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.1732}$
Table 32: Model performances on THCA
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $1500\scriptstyle\pm 0.1500 0.0$$200\scriptstyle\pm 0.0173 0.0$$348\scriptstyle\pm 0.0303 0.0$$0179\scriptstyle\pm 0.0155$ $0.1333\scriptstyle\pm 0.2309$ $0.1000\scriptstyle\pm 0.1732$
L3-FT(Med) $0439\scriptstyle\pm 0.0532 0.0$$167\scriptstyle\pm 0.0058 0.0$$200\scriptstyle\pm 0.0119 0.0$$101\scriptstyle\pm 0.0061 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(Med) + KG $0156\scriptstyle\pm 0.0145 0.0$$300\scriptstyle\pm 0.0265 0.0$$205\scriptstyle\pm 0.0187 0.0$$104\scriptstyle\pm 0.0095 0.0$$000\scriptstyle\pm 0.0000 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(QA) $0.5495\scriptstyle\pm 0.0165$ $0.5367\scriptstyle\pm 0.0153$ $0.5430\scriptstyle\pm 0.0158$ $0.3728\scriptstyle\pm 0.0148$ $0.8667\scriptstyle\pm 0.1155$ $0.8333\scriptstyle\pm 0.0577$
L3-FT(QA) + KG $0.5818\scriptstyle\pm 0.0490$ $0.5333\scriptstyle\pm 0.0416$ $0.5556\scriptstyle\pm 0.0365$ $0.3853\scriptstyle\pm 0.0353$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.1000$
G-Retriever $\mathbf{0.6237}\scriptstyle\pm\mathbf{0.0849}$ $0.3833\scriptstyle\pm 0.0231$ $0.4718\scriptstyle\pm 0.0097$ $0.3088\scriptstyle\pm 0.0083$ $0.5333\scriptstyle\pm 0.1155$ $0.7333\scriptstyle\pm 0.1155$
@@
Table 33: Model performances on BLCA
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0127\scriptstyle\pm 0.0111 0.0$$267\scriptstyle\pm 0.0231 0.0$$172\scriptstyle\pm 0.0149 0.0$$87\scriptstyle\pm 0.0076 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $0222\scriptstyle\pm 0.0385 0.0$$033\scriptstyle\pm 0.0058 0.0$$058\scriptstyle\pm 0.0100 0.0$$29\scriptstyle\pm 0.0051 0.000$$00\scriptstyle\pm 0.0000 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(QA) $0.5538\scriptstyle\pm 0.0498$ $0.5000\scriptstyle\pm 0.0200$ $0.5250\scriptstyle\pm 0.0301$ $0.3563\scriptstyle\pm 0.0279$ $0.7333\scriptstyle\pm 0.1155$ $0.8000\scriptstyle\pm 0.0000$
L3-FT(QA) + KG $\mathbf{0.5617}\scriptstyle\pm 0.0584$ $\mathbf{0.5367}\scriptstyle\pm 0.0231$ $\mathbf{0.5484}\scriptstyle\pm 0.0376$ $\mathbf{0.3784}\scriptstyle\pm 0.0356$ $0.8667\scriptstyle\pm 0.2309$ $0.8667\scriptstyle\pm 0.0577$
G-Retriever $0.5025\scriptstyle\pm 0.0481$ $0.4733\scriptstyle\pm 0.0058$ $0.4870\scriptstyle\pm 0.0253$ $0.3221\scriptstyle\pm 0.0223$ $0.6000\scriptstyle\pm 0.0000$ $0.6333\scriptstyle\pm 0.0577$
GALAX $0.5247\scriptstyle\pm 0.0081$ $0.5033\scriptstyle\pm 0.0289$ $0.5133\scriptstyle\pm 0.0115$ $0.3453\scriptstyle\pm 0.0104$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.1000}$
Table 34: Model performances on UCEC

### v2

@@ E.2 Human evaludation and LLM as judge
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0220\scriptstyle\pm 0.0249 0.0$$367\scriptstyle\pm 0.0404 0.0$$275\scriptstyle\pm 0.0308 0.0$$141\scriptstyle\pm 0.0159 0.0$$\scriptstyle\pm 0.0000 0.0000±0.00$$0\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $0375\scriptstyle\pm 0.0385 0.0$$067\scriptstyle\pm 0.0058 0.0$$111\scriptstyle\pm 0.0097 0.0$$056\scriptstyle\pm 0.0049 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics $0.3985\scriptstyle\pm 0.0669$ $0.3900\scriptstyle\pm 0.0721$ $0.3942\scriptstyle\pm 0.0695$ $0.2470\scriptstyle\pm 0.0530$ $0.9333\scriptstyle\pm 0.1155$ $0.8333\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics + KG $0.4200\scriptstyle\pm 0.0230$ $0.4100\scriptstyle\pm 0.0173$ $0.4149\scriptstyle\pm 0.0200$ $0.2619\scriptstyle\pm 0.0158$ $0.8667\scriptstyle\pm 0.1155$ $0.7000\scriptstyle\pm 0.1000$
G-Retriever + pre-GAT $0.3196\scriptstyle\pm 0.0307$ $0.2767\scriptstyle\pm 0.0058$ $0.2963\scriptstyle\pm 0.0168$ $0.1740\scriptstyle\pm 0.0115$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8000\scriptstyle\pm 0.0000$
RoG $0.4424\scriptstyle\pm 0.0239$ $0.4367\scriptstyle\pm 0.0289$ $0.4395\scriptstyle\pm 0.0264$ $0.2819\scriptstyle\pm 0.0219$ $0.8667\scriptstyle\pm 0.1155$ $0.7667\scriptstyle\pm 0.1155$
SubgraphRAG $0.4286\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.4242±0.0000$$42\scriptstyle\pm 0$.0000 $0.2692\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$ $0.7000\scriptstyle\pm 0.0000$
GNN-RAG $0.4433\scriptstyle\pm 0.0231$ $0.4433\scriptstyle\pm 0.0231$ $0.4433\scriptstyle\pm 0.0231$ $0.2850\scriptstyle\pm 0.0192$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.0000}$
GALAX $400\scriptstyle\pm 0.0000 0.440$$\scriptstyle\pm 0.0000 0.4400±0$$400\scriptstyle\pm 0.00$$0$ $0.2821\scriptstyle\pm 0.0000$ $0.8667\scriptstyle\pm 0.1155$ $0.6333\scriptstyle\pm 0.2309$
GALAX (Qwen2.5-7B) $\mathbf{0.4764}\scriptstyle\pm\mathbf{0.0055}$ $\mathbf{0.4700}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.4732}\scriptstyle\pm\mathbf{0.0027}$ $\mathbf{0.3099}\scriptstyle\pm\mathbf{0.0023}$ $0.8000\scriptstyle\pm 0.0000$ $\mathbf{0.9000}\scriptstyle\pm\mathbf{0.0000}$
Table 38: Model performance on $M$
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0275\scriptstyle\pm 0.0296 0.0$$167\scriptstyle\pm 0.0208 0.0$$156\scriptstyle\pm 0.0149 0.0$$079\scriptstyle\pm 0.0075 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(Med) + Omics + KG $1091\scriptstyle\pm 0.0630 0.0$$200\scriptstyle\pm 0.0000 0.0$$329\scriptstyle\pm 0.0027 0.0$$0167\scriptstyle\pm 0.0014$ $0.1333\scriptstyle\pm 0.1155$ $0.1333\scriptstyle\pm 0.1155$
L3-FT(CRISPR) + Omics $0.4001\scriptstyle\pm 0.0022$ $0.3800\scriptstyle\pm 0.0265$ $0.3894\scriptstyle\pm 0.0134$ $0.2418\scriptstyle\pm 0.0103$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics + KG $0.4753\scriptstyle\pm 0.0613$ $0.4367\scriptstyle\pm 0.0306$ $0.4549\scriptstyle\pm 0.0447$ $0.2952\scriptstyle\pm 0.0377$ $0.8000\scriptstyle\pm 0.0000$ $0.7333\scriptstyle\pm 0.0577$
G-Retriever + pre-GAT $0.4514\scriptstyle\pm 0.0890$ $0.3667\scriptstyle\pm 0.0808$ $0.4046\scriptstyle\pm 0.0850$ $0.2561\scriptstyle\pm 0.0690$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.0000$
RoG $0.4900\scriptstyle\pm 0.0615$ $0.3833\scriptstyle\pm 0.1550$ $0.4246\scriptstyle\pm 0.1182$ $0.2745\scriptstyle\pm 0.0997$ $0.9333\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.0577$
SubgraphRAG $0.4515\scriptstyle\pm 0.0122$ $0.3033\scriptstyle\pm 0.0289$ $0.3620\scriptstyle\pm 0.0175$ $0.2211\scriptstyle\pm 0.0130$ $0.9333\scriptstyle\pm 0.1155$ $0.8667\scriptstyle\pm 0.0577$
GNN-RAG $0.5166\scriptstyle\pm 0.0376$ $0.5133\scriptstyle\pm 0.0404$ $0.5150\scriptstyle\pm 0.0390$ $0.3474\scriptstyle\pm 0.0359$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8333\scriptstyle\pm 0.0577$
@@
GALAX (Qwen2.5-7B) $0.5000\scriptstyle\pm 0.0917$ $0.5000\scriptstyle\pm 0.0917$ $0.5000\scriptstyle\pm 0.0917$ $0.3367\scriptstyle\pm 0.0834$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.1000$
Table 39: Model performance on KIRC
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5458\scriptstyle\pm 0.0450$ $0.5333\scriptstyle\pm 0.0513$ $0.5394\scriptstyle\pm 0.0478$ $0.3703\scriptstyle\pm 0.0452$ $0.8667\scriptstyle\pm 0.2309$ $0.8667\scriptstyle\pm 0.1155$
L3-FT(CRISPR) + Omics + KG $0.4897\scriptstyle\pm 0.0182$ $0.4800\scriptstyle\pm 0.0265$ $0.4847\scriptstyle\pm 0.0219$ $0.3201\scriptstyle\pm 0.0192$ $0.8667\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.0000$
G-Retriever + pre-GAT $0.4160\scriptstyle\pm 0.0315$ $0.3733\scriptstyle\pm 0.0058$ $0.3933\scriptstyle\pm 0.0171$ $0.2449\scriptstyle\pm 0.0133$ $0.6000\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$
RoG $0.6207\scriptstyle\pm 0.0617$ $\mathbf{0.5867}\scriptstyle\pm\mathbf{0.0751}$ $\mathbf{0.6031}\scriptstyle\pm\mathbf{0.0689}$ $\mathbf{0.4340}\scriptstyle\pm\mathbf{0.0687}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
SubgraphRAG $\mathbf{0.6412}\scriptstyle\pm\mathbf{0.0261}$ $0.5300\scriptstyle\pm 0.1732$ $0.5714\scriptstyle\pm 0.1237$ $0.4067\scriptstyle\pm 0.1160$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9333\scriptstyle\pm 0.1155$
GNN-RAG $0.5898\scriptstyle\pm 0.0350$ $0.5733\scriptstyle\pm 0.0635$ $0.5812\scriptstyle\pm 0.0499$ $0.4108\scriptstyle\pm 0.0486$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
GALAX $0.5797\scriptstyle\pm 0.0251$ $0.5700\scriptstyle\pm 0.0265$ $0.5748\scriptstyle\pm 0.0254$ $0.4036\scriptstyle\pm 0.0247$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.1732$
GALAX (Qwen2.5-7B) $0.5786\scriptstyle\pm 0.1114$ $0.5700\scriptstyle\pm 0.1044$ $0.5742\scriptstyle\pm 0.1076$ $0.4079\scriptstyle\pm 0.1019$ $0.8667\scriptstyle\pm 0.1155$ $0.9333\scriptstyle\pm 0.0577$
Table 40: Model performance on THCA
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $1500\scriptstyle\pm 0.1500 0.0$$200\scriptstyle\pm 0.0173 0.0$$348\scriptstyle\pm 0.0303 0.0$$0179\scriptstyle\pm 0.0155$ $0.1333\scriptstyle\pm 0.2309$ $0.1000\scriptstyle\pm 0.1732$
L3-FT(Med) + Omics $0439\scriptstyle\pm 0.0532 0.0$$167\scriptstyle\pm 0.0058 0.0$$200\scriptstyle\pm 0.0119 0.0$$101\scriptstyle\pm 0.0061 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(Med) + Omics + KG $0156\scriptstyle\pm 0.0145 0.0$$300\scriptstyle\pm 0.0265 0.0$$205\scriptstyle\pm 0.0187 0.0$$104\scriptstyle\pm 0.0095 0.0$$000\scriptstyle\pm 0.0000 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics $0.5495\scriptstyle\pm 0.0165$ $0.5367\scriptstyle\pm 0.0153$ $0.5430\scriptstyle\pm 0.0158$ $0.3728\scriptstyle\pm 0.0148$ $0.8667\scriptstyle\pm 0.1155$ $0.8333\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics + KG $0.5818\scriptstyle\pm 0.0490$ $0.5333\scriptstyle\pm 0.0416$ $0.5556\scriptstyle\pm 0.0365$ $0.3853\scriptstyle\pm 0.0353$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.1000$
G-Retriever + pre-GAT $0.6237\scriptstyle\pm 0.0849$ $0.3833\scriptstyle\pm 0.0231$ $0.4718\scriptstyle\pm 0.0097$ $0.3088\scriptstyle\pm 0.0083$ $0.5333\scriptstyle\pm 0.1155$ $0.7333\scriptstyle\pm 0.1155$
RoG $0.5696\scriptstyle\pm 0.0350$ $0.5367\scriptstyle\pm 0.0635$ $0.5524\scriptstyle\pm 0.0499$ $0.3827\scriptstyle\pm 0.0486$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
SubgraphRAG $0.5495\scriptstyle\pm 0.0000$ $0.5000\scriptstyle\pm 0.0000$ $0.5236\scriptstyle\pm 0.0000$ $0.3546\scriptstyle\pm 0.0000$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
GNN-RAG $\mathbf{0.6465}\scriptstyle\pm\mathbf{0.0350}$ $0.4400\scriptstyle\pm 0.1386$ $0.5127\scriptstyle\pm 0.0782$ $0.3473\scriptstyle\pm 0.0731$ $0.9333\scriptstyle\pm 0.1155$ $0.8333\scriptstyle\pm 0.0577$
@@
GALAX (Qwen2.5-7B) $0.5790\scriptstyle\pm 0.0350$ $0.4967\scriptstyle\pm 0.0751$ $0.5312\scriptstyle\pm 0.0356$ $0.3622\scriptstyle\pm 0.0333$ $0.8000\scriptstyle\pm 0.2000$ $0.8333\scriptstyle\pm 0.0577$
Table 41: Model performance on BLCA
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0127\scriptstyle\pm 0.0111 0.0$$267\scriptstyle\pm 0.0231 0.0$$172\scriptstyle\pm 0.0149 0.0$$87\scriptstyle\pm 0.0076 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $0222\scriptstyle\pm 0.0385 0.0$$033\scriptstyle\pm 0.0058 0.0$$058\scriptstyle\pm 0.0100 0.0$$29\scriptstyle\pm 0.0051 0.000$$00\scriptstyle\pm 0.0000 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics $0.5538\scriptstyle\pm 0.0498$ $0.5000\scriptstyle\pm 0.0200$ $0.5250\scriptstyle\pm 0.0301$ $0.3563\scriptstyle\pm 0.0279$ $0.7333\scriptstyle\pm 0.1155$ $0.8000\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics + KG $0.5617\scriptstyle\pm 0.0584$ $0.5367\scriptstyle\pm 0.0231$ $0.5484\scriptstyle\pm 0.0376$ $0.3784\scriptstyle\pm 0.0356$ $0.8667\scriptstyle\pm 0.2309$ $0.8667\scriptstyle\pm 0.0577$
G-Retriever + pre-GAT $0.5025\scriptstyle\pm 0.0481$ $0.4733\scriptstyle\pm 0.0058$ $0.4870\scriptstyle\pm 0.0253$ $0.3221\scriptstyle\pm 0.0223$ $0.6000\scriptstyle\pm 0.0000$ $0.6333\scriptstyle\pm 0.0577$
RoG $0.5633\scriptstyle\pm 0.0231$ $0.5633\scriptstyle\pm 0.0231$ $0.5633\scriptstyle\pm 0.0231$ $0.3924\scriptstyle\pm 0.0226$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0577}$
SubgraphRAG $0.5457\scriptstyle\pm 0.0075$ $0.4633\scriptstyle\pm 0.1501$ $0.4922\scriptstyle\pm 0.1001$ $0.3302\scriptstyle\pm 0.0851$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.1732$
GNN-RAG $0.5270\scriptstyle\pm 0.1006$ $0.4900\scriptstyle\pm 0.1646$ $0.5056\scriptstyle\pm 0.1376$ $0.3456\scriptstyle\pm 0.1177$ $0.6667\scriptstyle\pm 0.5774$ $0.6000\scriptstyle\pm 0.5196$
GALAX $0.5247\scriptstyle\pm 0.0081$ $0.5033\scriptstyle\pm 0.0289$ $0.5133\scriptstyle\pm 0.0115$ $0.3453\scriptstyle\pm 0.0104$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.1000$
GALAX (Qwen2.5-7B) $\mathbf{0.5852}\scriptstyle\pm\mathbf{0.0091}$ $\mathbf{0.5733}\scriptstyle\pm\mathbf{0.0115}$ $\mathbf{0.5791}\scriptstyle\pm\mathbf{0.0015}$ $\mathbf{0.4076}\scriptstyle\pm\mathbf{0.0015}$ $0.9333\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.0000$
Table 42: Model performance on UCEC